# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v3

@@ C.3 Full results partitioned by model capabilities
    Lag-Llama 0.329 $\pm$ 0.004 29.074 $\pm$ 0.783 0.355 $\pm$ 0.007 0.181 $\pm$ 0.003 0.324 $\pm$ 0.003 0.272 $\pm$ 0.006 0.342 $\pm$ 0.006 0.386 $\pm$ 0.009 0.449 $\pm$ 0.012
    Chronos-Tiny 0.328 $\pm$ 0.001 26.277 $\pm$ 0.416 0.400 $\pm$ 0.002 0.144 $\pm$ 0.003 0.297 $\pm$ 0.002 0.229 $\pm$ 0.002 0.290 $\pm$ 0.005 0.382 $\pm$ 0.002 0.440 $\pm$ 0.003
    Chronos-Mini 0.341 $\pm$ 0.001 27.560 $\pm$ 0.394 0.412 $\pm$ 0.002 0.147 $\pm$ 0.002 0.302 $\pm$ 0.002 0.248 $\pm$ 0.002 0.305 $\pm$ 0.004 0.378 $\pm$ 0.003 0.436 $\pm$ 0.004
    Chronos-Small 0.328 $\pm$ 0.002 25.217 $\pm$ 0.349 0.388 $\pm$ 0.003 0.144 $\pm$ 0.002 0.287 $\pm$ 0.002 0.248 $\pm$ 0.003 0.290 $\pm$ 0.003 0.358 $\pm$ 0.005 0.412 $\pm$ 0.006
    Chronos-Base 0.672 $\pm$ 0.003 29.202 $\pm$ 0.353 0.702 $\pm$ 0.002 0.143 $\pm$ 0.002 1.023 $\pm$ 0.006 0.261 $\pm$ 0.003 0.299 $\pm$ 0.004 1.643 $\pm$ 0.009 2.187 $\pm$ 0.012
    Chronos-Large 0.326 $\pm$ 0.002 24.459 $\pm$ 0.380 0.385 $\pm$ 0.002 0.138 $\pm$ 0.002 0.288 $\pm$ 0.002 0.249 $\pm$ 0.002 0.295 $\pm$ 0.003 0.362 $\pm$ 0.003 0.417 $\pm$ 0.004
    TimeGEN 0.354 $\pm$ 0.000 34.432 $\pm$ 0.165 0.402 $\pm$ 0.000 0.176 $\pm$ 0.000 0.308 $\pm$ 0.000 0.279 $\pm$ 0.000 0.324 $\pm$ 0.000 0.377 $\pm$ 0.000 0.431 $\pm$ 0.000
    Moirai-small 0.565 $\pm$ 0.031 33.878 $\pm$ 0.421 0.429 $\pm$ 0.005 0.671 $\pm$ 0.146 0.468 $\pm$ 0.076 0.566 $\pm$ 0.017 1.204 $\pm$ 0.271 0.376 $\pm$ 0.009 0.426 $\pm$ 0.012
    Moirai-base 0.624 $\pm$ 0.013 33.138 $\pm$ 0.378 0.410 $\pm$ 0.006 0.600 $\pm$ 0.053 0.680 $\pm$ 0.028 0.690 $\pm$ 0.019 1.147 $\pm$ 0.099 0.375 $\pm$ 0.010 0.441 $\pm$ 0.013
    Moirai-large 0.520 $\pm$ 0.006 27.016 $\pm$ 0.843 0.414 $\pm$ 0.004 0.155 $\pm$ 0.004 0.260 $\pm$ 0.003 0.751 $\pm$ 0.015 0.276 $\pm$ 0.008 0.337 $\pm$ 0.007 0.397 $\pm$ 0.010
@@ C.5 Significant failures per model
     UniTime 0
     Time-LLM (ETTh1) 2
@@ C.6 Testing the Statistical Significance of the Relevance of Context
0.378 ± 0.001 0.364 ± 0.001 1.000
@@
0.413 ± 0.009 0.432 ± 0.010 0.294
direct prompt
Mixtral-8x7B-Inst
0.414 ± 0.007 0.699 ± 0.006 1.89 $\times$ 10-${}^{\text{-6}}$
@@
0.480 ± 0.016 0.406 ± 0.006 0.973
@@ C.10 The effect of significant failures on the aggregate performance of models
Figure 16: Histogram of the RCPRS (lower is better) of the direct prompt  - Mixtral-8x7B-Instruct model on each task, with and without context (with the weighting scheme detailed in Sec. A.5). With context, the model gets an RCRPS close to zero in a large number of tasks (also achieving a high win rate as seen in Figure 5). However, there is also a long tail of tasks with high RCRPS values with context, dominating and worsening the model’s aggregate RCRPS.
@@ D.1.3 Example Prompt
Scenario: Suppose that there is an accident on the road and there is 40.0% of the usual traffic from 2024-04-24 17:00:00 for 6 hours.
@@ D.2.1 Method
Suppose that there is an accident on the road and there is 40.0% of the usual traffic from 2024-04-24 17:00:00 for 6 hours.
@@ D.2.2 Implementation Details
D.3 UniTime and Time-LLM
@@ E.5 Covariance of two CRPS estimators
To get an estimate of the covariance using our $M$ samples, we can estimate each of these terms using their respective unbiased estimators. Once we have computed an estimate of the variance for a single task instance, the overall variance for a full task is computed using the formula for the variance of the average of multiple independent variables. One slight disadvantage of using this method, is that it offers no guarantee that the RCPRS variance estimate will be non-negative, so in the rare cases where the estimate for the variance of a full task is negative, we clip it to 0.

### v4

@@ C.3 Full results partitioned by model capabilities
    Lag-Llama 0.327 $\pm$ 0.004 30.451 $\pm$ 0.819 0.353 $\pm$ 0.007 0.181 $\pm$ 0.003 0.324 $\pm$ 0.003 0.269 $\pm$ 0.006 0.342 $\pm$ 0.006 0.386 $\pm$ 0.009 0.449 $\pm$ 0.012
    Chronos-Tiny 0.328 $\pm$ 0.001 27.487 $\pm$ 0.441 0.400 $\pm$ 0.002 0.144 $\pm$ 0.003 0.297 $\pm$ 0.002 0.229 $\pm$ 0.002 0.290 $\pm$ 0.005 0.382 $\pm$ 0.002 0.440 $\pm$ 0.003
    Chronos-Mini 0.341 $\pm$ 0.001 28.893 $\pm$ 0.428 0.412 $\pm$ 0.002 0.147 $\pm$ 0.002 0.302 $\pm$ 0.002 0.248 $\pm$ 0.002 0.305 $\pm$ 0.004 0.378 $\pm$ 0.003 0.436 $\pm$ 0.004
    Chronos-Small 0.328 $\pm$ 0.002 26.524 $\pm$ 0.372 0.388 $\pm$ 0.003 0.144 $\pm$ 0.002 0.287 $\pm$ 0.002 0.248 $\pm$ 0.003 0.290 $\pm$ 0.003 0.358 $\pm$ 0.005 0.412 $\pm$ 0.006
    Chronos-Base 0.672 $\pm$ 0.003 30.601 $\pm$ 0.375 0.702 $\pm$ 0.002 0.143 $\pm$ 0.002 1.023 $\pm$ 0.006 0.261 $\pm$ 0.003 0.299 $\pm$ 0.004 1.643 $\pm$ 0.009 2.187 $\pm$ 0.012
    Chronos-Large 0.326 $\pm$ 0.002 25.602 $\pm$ 0.399 0.385 $\pm$ 0.002 0.138 $\pm$ 0.002 0.288 $\pm$ 0.002 0.249 $\pm$ 0.002 0.295 $\pm$ 0.003 0.362 $\pm$ 0.003 0.417 $\pm$ 0.004
    TimeGEN 0.353 $\pm$ 0.000 35.924 $\pm$ 0.167 0.401 $\pm$ 0.000 0.176 $\pm$ 0.000 0.308 $\pm$ 0.000 0.278 $\pm$ 0.000 0.324 $\pm$ 0.000 0.377 $\pm$ 0.000 0.431 $\pm$ 0.000
    Moirai-small 0.565 $\pm$ 0.031 36.038 $\pm$ 0.438 0.429 $\pm$ 0.005 0.671 $\pm$ 0.146 0.468 $\pm$ 0.076 0.566 $\pm$ 0.017 1.204 $\pm$ 0.271 0.376 $\pm$ 0.009 0.426 $\pm$ 0.012
    Moirai-base 0.624 $\pm$ 0.013 35.263 $\pm$ 0.407 0.410 $\pm$ 0.006 0.600 $\pm$ 0.053 0.680 $\pm$ 0.028 0.690 $\pm$ 0.019 1.147 $\pm$ 0.099 0.375 $\pm$ 0.010 0.441 $\pm$ 0.013
    Moirai-large 0.520 $\pm$ 0.006 28.635 $\pm$ 0.862 0.414 $\pm$ 0.004 0.155 $\pm$ 0.004 0.260 $\pm$ 0.003 0.751 $\pm$ 0.015 0.276 $\pm$ 0.008 0.337 $\pm$ 0.007 0.397 $\pm$ 0.010
@@ C.5 Significant failures per model
UniTime
0
Time-LLM (ETTh1)
2
ChatTime-Base
21
ChatTime-Chat
23
@@ C.6 Testing the Statistical Significance of the Relevance of Context
0.378 ± 0.001 0.364 ± 0.001 1 - 8.08 $\times$ 10-${}^{\text{-7}}$
direct prompt
Mixtral-8x7B-Inst
0.413 ± 0.007 0.699 ± 0.006 1.88 $\times$ 10-${}^{\text{-6}}$
@@
0.413 ± 0.009 0.432 ± 0.010 0.287
@@
0.481 ± 0.016 0.406 ± 0.006 0.975
ChatTime-Chat
0.557 ± 0.001 0.554 ± 0.001 0.086
ChatTime-Base
0.568 ± 0.001 0.556 ± 0.002 1 - 1.35 $\times$ 10-${}^{\text{-4}}$
@@ C.10 The effect of significant failures on the aggregate performance of models
Figure 16: Histogram of the RCPRS (lower is better) of the direct prompt  - Qwen-2.5-0.5B-Instruct model on each task, with and without context. With context, the RCRPS is close to zero for a large number of tasks, but there is also a long tail of tasks with high RCRPS values, dominating and worsening the model’s aggregate RCRPS.
@@ D.1.3 Example Prompt
Scenario: Suppose that there is an accident on the road and there is 40.0%
@@ D.2.1 Method
Suppose that there is an accident on the road and there is 40.0%
@@ D.2.2 Implementation Details
D.3 ChatTime
We evaluate the released ChatTime-Base (https://huggingface.co/ChengsenWang/ChatTime-1-7B-Base) and ChatTime-Chat (https://huggingface.co/ChengsenWang/ChatTime-1-7B-Chat) models zero-shot, as per the instructions in the authors’ GitHub repository (https://github.com/ForestsKing/ChatTime).
D.4 UniTime and Time-LLM
@@ E.5 Covariance of two CRPS estimators
To get an estimate of the covariance using our $M$ samples, we can estimate each of these terms using their respective unbiased estimators. Once we have computed an estimate of the variance for a single task instance, the overall variance for a full task is computed using the formula for the variance of the average of multiple independent variables. One slight disadvantage of using this method is that it offers no guarantee that the RCPRS variance estimate will be non-negative, so in the rare cases where the estimate for the variance of a full task is negative, we clip it to 0.
E.6 Comparison of Statistical Properties of Various Scoring Rules
Table 12: Comparison of Statistical Properties of Various Scoring Rules. The * indicates that, to be proper, it would need that different seeds are independent, which cannot be guaranteed by CiK, but could happen in other applications.
Metric Proper Scoring Rule Domain Invariance
Additive Multiplicative
Brier Score Yes Discrete Yes Yes
CRPS Yes Continuous Yes No
TwCRPS Yes Continuous Yes No
CRPS skill score No Continuous Yes Yes
MAV-Scaled CRPS No Continuous No Yes
RCRPS No* Continuous Yes Yes
Tab. 12 describes a few statistical properties for both commonly used scoring rules and our RCPRS. For the invariance (additive and multiplicative) properties, we indicate whether the scoring rule remains unchanged if all relevant quantities (forecast, ground truth, threshold, and constraint parameters) are modified by adding a constant to them or by multiplying them by a constant. By MAV-Scaled CRPS, we denote the common approach in the forecasting literature to normalize the CRPS by dividing it by the Mean Absolute Values of the ground-truth, instead of reporting the original CRPS values.